### v4

# Quantum advantage from negativity of work quasiprobability distributions

Abstract
Quantum batteries can be charged by performing a work “instantaneously” in the limit of large number of cells. In general, the work exhibits statistics that can be represented by a quasiprobability in the presence of initial quantum coherence in the energy basis. Here we show that these two concepts of quantum thermodynamics, which apparently appear disconnected, show a simple relation. Specifically, if the work distribution shows negativity in some time interval, then we can surely get quantum advantage in the charging process.
I Introduction
Quantum batteries Campaioli19 ; Bhattacharjee21 ; Campaioli24 are thought of as quantum systems capable of storing energy. This energy can be used to perform useful work, e.g., through a cyclic change of some parameters of the quantum system Alicki13 . The maximum work extractable is called ergotropy Allahverdyan04 and can be related to some quantum features of the initial state, e.g., quantum correlations Francica17 ; Bernards19 ; Touil22 ; Cruz22 ; Francica22 and quantum coherence Francica20 ; Cakmak20 ; Niu24 . This amount of work can also be extracted without generating entanglement during the process Hovhannisyan13 . Furthermore, all operations that do not generate this useful work have recently been investigated Francica25r . Concerning the charging of the battery, when the battery is formed by $N$ cells (or copies) the duration time tends to zero as $N$ increases, and thus a quantum advantage is achieved Binder15 ; Campaioli17 ; Gyhm22 . Among the several proposals for the realization of quantum batteries (by using many-body interactions in spin systems Julia-Farre20 , with a cavity assisted charging Ferraro18 ; Andolina19 ; Andolina18 ; Farina19 ; Zhang19 ; Andolina192 ; Crescente20 , in disordered chains Rossini19 and fermionic Sachdev-Ye-Kitaev (SYK) interactions Rossini20 ; Rosa20 ; Carrega21 ; Gyhm24 , to name a few), rescaled SYK interactions Francica25 , where the (correlated) disorder is chosen in order to saturate the Bhatia-Davis bound Bhatia00 , allow to achieve an optimal and extensive quantum advantage.
On the other hand, the work done exhibits a statistics that satisfies fluctuation theorems Jarzynski97 ; Crooks99 when it is taken into account with a two-point-measurement scheme Kurchan00 ; Tasaki00 ; Talkner07 ; Campisi11 , for an incoherent initial state. If there is initial quantum coherence with respect to the energy basis, a no-go theorem Perarnau-Llobet17 indicates that the statistics can be represented with some quasiprobability distribution Baumer18 ; Gherardini24 ; Arvidsson-Shukur24 , i.e., a normalized distribution that also takes negative Allahverdyan14 ; Solinas15 or complex Lostaglio24 values. This can be seen as a consequence of the presence of quantum contextuality Lostaglio18 . Concerning the form of the quasiprobability, it can be derived by starting from some axioms, e.g., the first law of thermodynamics. For weak conditions, no unique quasiprobability distribution is selected, but we get a class of quasiprobability distributions Francica22a ; Francica22b . A description in terms of quasiprobabilities allows one to perform an optimization with a utility function Francica24a ; Francica24b , i.e., based on higher moments of the work, and fluctuation theorems can also be formulated Francica24ft . However, it is not clear what the role of quasiprobability of work in the charging quantum advantage is, where the duration time plays a key role. Here, we try to answer this question.
To do this, we start by introducing the charging process in Sec. II and the work quasiprobability distributions in Sec. III. In particular, we give an intuitive simple calculation in Sec. III.1. Then, we explain how the negativity of the work quasiprobability distribution gives a sufficient condition for achieving quantum advantage in Sec. IV. We further investigate this result with the help of a fermionic model in Sec. V. In the end, we summarize and discuss possible applications of our result in Sec. VI.
II Charging process
In order to introduce the charging process, we focus on a quantum battery having $N$ cells and the free Hamiltonian $H_{0}=\sum_{i=1}^{N}h_{i}$, where $h_{i}$ is the local Hamiltonian of the i-th cell. In our process, the battery is initially prepared in the ground-state of $H_{0}$, which is $|E_{0}\rangle=|0\rangle^{\otimes N}$ with $|0\rangle$ such that $h_{i}|0\rangle=0$ for all $i=1,\ldots,N$. Thus, the lowest energy of the Hamiltonian $H_{0}$ is $E^{0}_{min}=E_{0}=0$. The charging process leads the system to the final state $|\psi(\tau)\rangle=U_{\tau,0}|\psi(0)\rangle$ where $U_{t,0}$ is the time-evolution unitary operator generated by the time-dependent Hamiltonian $H(t)$ with $t\in[0,\tau]$ such that $H(0)=H(\tau)=H_{0}$. For simplicity, let us focus on a time-dependent Hamiltonian such that $H(t)=H_{0}$ for $t=0,\tau$ and $H(t)=H_{1}$ for $t\in(0,\tau)$. This situation can be realized by performing two sudden quenches at the initial ($t=0$) and final ($t=\tau$) times. In this case, the time-evolution unitary operator is $U_{t,0}=e^{-iH_{1}t}$ when $t\in(0,\tau)$. We consider a Hamiltonian $H_{1}$ such that $H_{1},H_{0}]\neq 0$ in order to achieve a non-trivial dynamics with a final state $|\psi(\tau)\rangle\propto|E_{1}\rangle$, where $|E_{1}\rangle$ is the highest excited state of $H_{0}$, with energy $E_{1}=E^{0}_{max}\sim N$, since we have $|E_{1}\rangle=|1\rangle^{\otimes N}$ where $|1\rangle$ is the eigenstate of $h_{i}$ with maximum energy. Thus, at the end of the process we get a fully charged battery. In order to optimize this charging process, we want to minimize the duration time $\tau$ of the process. To do this, the choice of the charging Hamiltonian $H_{1}$ is typically constrained by some conditions, e.g., the gap $\Delta E_{1}=E_{max}-E_{min}$ between the maximum and minimum eigenvalues of $H_{1}$, which are $E_{max}$ and $E_{min}$, is not larger than the gap $\Delta E_{0}=E^{0}_{max}$ between the maximum and minimum eigenvalues of $H_{0}$, which are $E^{0}_{max}\sim N$ and $E^{0}_{min}=0$. If no interactions among the cells are allowed (parallel charging), the charging Hamiltonian has the form $H_{1}=\sum_{i}v_{i}$, where $v_{i}$ is a local operator of the i-th cell, and we get a minimum duration time $\tau$ that remains constant and nonzero as $N\to\infty$. On the other hand, if interactions are allowed, e.g., terms like $\sum_{i_{1},i_{2},\ldots,i_{q}}v_{i_{1}}\otimes v_{i_{2}}\otimes\cdots\otimes v_{i_{q}}$ may be present in the Hamiltonian $H_{1}$, we can outperform the parallel charging case, getting a duration time $\tau\to 0$ as $N\to\infty$, and thus we achieve a charging quantum advantage. For instance, we can get $\tau\sim N^{-z}$ for some exponent $z>0$.
II.1 Sufficient condition for quantum advantage
In order to achieve a sufficient condition to get $\tau\to 0$ as $N\to\infty$, we note that the average work done in the process in general is $W=\langle H_{0}\rangle_{\tau}-\langle H_{0}\rangle_{0}$, where in general we defined $\langle X\rangle_{t}=\text{Tr}\left\{X\rho(t)\right\}$ for any operator $X$ and a density matrix $\rho(t)$. In our case, $t\in[0,\tau]$ and $\rho(t)$ is the time-evolved pure state $\rho(t)=|\psi(t)\rangle\langle\psi(t)|$ with $|\psi(t)\rangle=U_{t,0}|\psi(0)\rangle$, so that $\langle X\rangle_{t}=\langle\psi(t)|X|\psi(t)\rangle$. In particular, we have $\langle H_{0}\rangle_{t}=\sum_{i}\langle e^{iH_{1}t}h_{i}e^{-iH_{1}t}\rangle_{0}$ which can be expanded in a power series in $t$ by using the Baker-Campbell-Hausdorff formula, i.e., $\langle H_{0}\rangle_{t}=\sum_{i}\langle H_{1}h_{i}H_{1}\rangle_{0}t^{2}+\cdots$, where we noted that $|\psi(0)\rangle=|E_{0}\rangle$ and $E_{0}=E^{0}_{min}=0$. Since $|\psi(\tau)\rangle\propto|E_{1}\rangle$ for our process, we get
$\frac{E_{1}}{N}=\frac{1}{N}\sum_{i}\langle H_{1}h_{i}H_{1}\rangle_{0}{\tau}^{2}+\cdots=\sum_{k}\sum_{i}\frac{w_{ik}}{N}{\tau}^{k}\,$, (1)
where we note that the series is convergent. To see how $\tau$ scales with $N$, we can use Eq. (1), which of course is satisfied for some duration time $\tau$. Thus, if for some $k$ the coefficient of the k-th term $\sum_{i}w_{ik}{\tau}^{k}/N$ is unbounded, i.e.,
$\left|\sum_{i}\frac{w_{ik}}{N}\right|\to\infty$ (2)
as $N\to\infty$, then $\tau\to 0$ since the series in Eq. (1) converges to the sum $E_{1}/N\sim{\it O}(1)$, so that the series is bounded as $N\to\infty$. We get the sufficient condition $|\sum_{i}w_{i2}|/N=|\langle H_{1}H_{0}H_{1}\rangle_{0}|/N\to\infty$ that implies $\tau\to 0$. Thus, in this case, there is quantum advantage. In particular, when $w_{ik}\sim N^{kz}$ for some $z>0$ and for $N^{\prime}\sim N$ indices $i$, we deduce that $\tau\sim N^{-z}$.
III Quasiprobability distribution
Having introduced our charging process, let us focus on the work done in a time interval $t_{1},t_{2}]$ for a generic out-of-equilibrium process generated by a time-dependent Hamiltonian $H(t)$. We are considering a time interval $t_{1},t_{2}]$ different from the one previously considered, which was $0,\tau]$, since we are going to focus on $t_{1}\in[0,t_{2})$ and $t_{2}\in(0,\tau]$ in our charging process, later in the paper. In general, the statistics of the work can be represented by some quasiprobability distribution, and for our purposes we focus on
$p_{q}(w)=\sum_{i,j,k}\text{Re}\langle E_{i}^{t_{1}}|\rho(t_{1})|E_{j}^{t_{1}}\rangle\langle E_{j}^{t_{1}}|U^{\dagger}_{t_{2},t_{1}}|E_{k}^{t_{2}}\rangle$
$\times\langle E_{k}^{t_{2}}|U_{t_{2},t_{1}}|E_{i}^{t_{1}}\rangle\delta(w-E_{k}^{t_{2}}+qE_{i}^{t_{1}}+(1-q)E_{j}^{t_{1}})\,$, (3)
where $q$ is a real number, $\rho(t_{1})$ is the density matrix at the time $t=t_{1}$, and the instantaneous eigenvalues $E_{i}^{t}$ and eigenstates $|E_{i}^{t}\rangle$ are defined through the spectral decomposition $H(t)=\sum_{i}E_{i}^{t}|E_{i}^{t}\rangle\langle E_{i}^{t}|$. We recall that the time-evolution operator is obtained as $U_{t_{2},t_{1}}=\mathcal{T}e^{-i\int_{t_{1}}^{t_{2}}H(t)dt}$, where $\mathcal{T}$ is the time ordering operator. Thus, the moments of the work are defined as $\langle w^{n}\rangle=\int w^{n}p_{q}(w)dw$. In particular, the first three moments read $\langle w\rangle=\langle(H^{(H)}(t_{2})-H(t_{1}))\rangle_{t_{1}}$, $\langle w^{2}\rangle=\langle(H^{(H)}(t_{2})-H(t_{1}))^{2}\rangle_{t_{1}}$ and
$\langle w^{3}\rangle$ $=$ $\langle(H^{(H)}(t_{2})-H(t_{1}))^{3}\rangle_{t_{1}}$ (4)
$+\frac{1}{2}\langle[H(t_{1})+H^{(H)}(t_{2}),[H(t_{1}),H^{(H)}(t_{2})]]\rangle_{t_{1}}$
$-3q(1-q)\langle[H(t_{1}),[H(t_{1}),H^{(H)}(t_{2})]]\rangle_{t_{1}}\,$,
where the average is calculated with respect to $\rho(t_{1})$ as $\langle X\rangle_{t}=\text{Tr}\left\{X\rho(t)\right\}$ for any operator $X$ and we defined $H^{(H)}(t_{2})=U^{\dagger}_{t_{2},t_{1}}H(t_{2})U_{t_{2},t_{1}}$. We note that only the first two moments do not depend on the parameter $q$, whereas the higher moments, which will play a key role in our analysis, depend on the particular representation chosen, i.e., on the value of $q$. This quasiprobability distribution $p_{q}(w)$ satisfies some conditions Francica22b and helps us to understand what work statistics in quantum regime is. When the quasiprobability distribution $p_{q}(w)$ also takes negative values, we can get some advantage. As noted in Ref. Francica24b , when there is negativity the moments are not constrained by the Jensen’s inequality. Then, given a strictly increasing function $f(w)$ that is concave, i.e., such that $f^{\prime}(w)>0$ and $f^{\prime\prime}(w)\leq 0$, if $p_{q}(w)\geq 0$ for all $w$, from the Jensen’s inequality we get the lower bound
$\langle w\rangle\geq f^{-1}(\langle f(w)\rangle)\,.$ (5)
If $f(w)=aw+b$, with $a>0$, this bound is trivially saturated, but for a function $f(w)$ such that $\min_{w}-f^{\prime\prime}(w)/f^{\prime}(w)\geq\varepsilon$ for a given $\varepsilon>0$, the bound cannot be saturated and if the bound is violated then there is negativity. We can refer to $f(w)$ as utility function, as usual. Thus, when $\langle w\rangle$ is negative for some processes of work extraction, we can get a larger amount of work extracted $w_{ex}=-\langle w\rangle$ going beyond the bound given by the utility $f(w)$, provided that these processes show quasiprobability distributions of work also taking negative values. However, differently from the case discussed, it is not clear how the negativity of the quasiprobability can be related to the charging quantum advantage, which involves the duration time $\tau$ and the work moments seem to play no role.
Negativity can be measured with $\mathcal{N}=\int|p_{q}(w)|dw$, which is larger than 1 if and only if $p_{q}(w)$ takes also negative values Francica23 . We note that $|\chi_{q}(u)|\leq\mathcal{N}$, where we defined the characteristic function $\chi_{q}(u)=\langle e^{iwu}\rangle$. Thus, if we get $|\chi_{q}(u)|>1$ then there is negativity. Thus, negativity can be inferred experimentally by measuring the characteristic function $\chi_{q}(u)$ via an interferometric measurement scheme (see, e.g., Ref. Francica22a ).
III.1 Quantum advantage and work quasiprobability
To explain the intuition behind our main result, we focus on the charging process of Sec. II with $H_{1}=N\lambda(|E_{1}\rangle\langle E_{0}|+|E_{0}\rangle\langle E_{1}|)$. In this case, the time-evolved state is $|\psi(t)\rangle=\cos(\omega t/2)|E_{0}\rangle-i\sin(\omega t/2)|E_{1}\rangle$, where $\omega=2N\lambda$. The duration time is $\tau=\pi/\omega$, so that we get the final state $|\psi(\tau)\rangle=-i|E_{1}\rangle$ and the battery is fully charged. Let us examine the statistics of the work done in a time interval $t_{1},t_{2}]$. For $t_{1}=0$ there is no coherence with respect to the energy basis at the initial time $t_{1}$, then the work distribution is a probability distribution. For $t_{1}\in(0,t_{2})$ and $t_{2}<\tau$ we get $H(t_{1})=H(t_{2})=H_{1}$ and since the time-evolution is generated by $H_{1}$, we get a trivial situation with a probability distribution for the work. In contrast, the case $t_{1}\in(0,\tau)$ and $t_{2}=\tau$ appears us to be more interesting. In particular for $t_{1}=\tau/2$ and $t_{2}=\tau$ it is easy to see that the work quasiprobability distribution $p_{q}(w)$ reads
$p_{q}(w)=\frac{1}{4}\sum_{i=\pm,k=0,1}\delta(w-E_{k}+E_{i})+\delta p_{q}(w)\,$, (6)
where $E_{\pm}$ are the only eigenvalues of $H_{1}$ different from zero and $|E_{\pm}\rangle$ are the corresponding eigenstates. In detail, we defined
$\delta p_{q}(w)$ $=$ $\frac{1}{4}\sum_{k=0,1}(-1)^{k+1}(\delta(w-E_{k}+qE_{+}+(1-q)E_{-})$ (7)
$+\delta(w-E_{k}+qE_{-}+(1-q)E_{+}))\,$,
which of course also takes negative values. Then, the quasiprobability distribution $p_{q}(w)$ is a probability distribution for $q=0,1$, explicitly $p_{0}(w)=\frac{1}{2}\delta(w-E_{1}+E_{-})+\frac{1}{2}\delta(w-E_{1}+E_{+})$. In contrast, for $q\neq 0,1$, $p_{q}(w)$ also takes negative values due to the term $\delta p_{q}(w)$. Thus, the charging quantum advantage appears to be related to the negativity of $p_{q}(w)$.
III.2 Leggett-Garg inequality
Let us focus on the work $w$ done in the time interval $t_{1},\tau]$ with $t_{1}\in(0,\tau)$. We can consider the works $w_{\tau}$ and $w_{t_{1}}$ done in the time intervals $0,\tau]$ and $0,t_{1}]$. Since $\rho(0)$ is incoherent with respect to the energy basis, these works have the probability distribution $p(w_{t})=\sum_{i,k}p_{i}p_{k|i}\delta(w_{t}-E^{t}_{k}+E^{0}_{i})$ with $t=\tau,t_{1}$, where $p_{i}=\langle E^{0}_{i}|\rho(0)|E^{0}_{i}\rangle$ and $p_{k|i}=|\langle E^{t}_{k}|U_{t,0}|E^{0}_{i}\rangle|^{2}$. Thus, if $p_{q}(w)<0$ for some $w$, then there is no joint probability distribution $p(w_{\tau},w_{t_{1}})$, since $w=w_{\tau}-w_{t_{1}}$ and thus $p(w_{\tau},w_{t_{1}})$ is such that $\int\delta(w-w_{\tau}+w_{t_{1}})p(w_{\tau},w_{t_{1}})dw_{\tau}dw_{t_{1}}=p_{q}(w)$. Furthermore, if there exists a joint probability distribution $p(w_{\tau},w_{t_{1}})$, we get the inequality for the covariance
$|\langle w_{\tau}w_{t_{1}}\rangle-\langle w_{\tau}\rangle\langle w_{t_{1}}\rangle|\leq\sigma^{2}_{\tau}\sigma^{2}_{t_{1}}\,$, (8)
where $\sigma^{2}_{t}=\langle w_{t}^{2}\rangle-\langle w_{t}\rangle^{2}$. This inequality can be expressed in terms of the variance $\sigma^{2}=\langle w^{2}\rangle-\langle w\rangle^{2}$ as
$|\sigma^{2}_{\tau}+\sigma^{2}_{t_{1}}-\sigma^{2}|\leq 2\sigma^{2}_{\tau}\sigma^{2}_{t_{1}}\,$, (9)
which recalls a Leggett-Garg inequality Leggett85 ; Miller18 . In particular, Eq. (9) does not depend on $q$ and thus on the quasiprobability representation chosen. Therefore, if the inequality in Eq. (9) is violated, then there is no joint probability $p(w_{\tau},w_{t_{1}})$ and we can get a distribution $p_{q}(w)$ that is negative for some $w$. It is easy to check that for the charging process under consideration the inequality is not violated, since $\sigma^{2}_{\tau}=0$ and $\sigma^{2}_{t_{1}}=\sigma^{2}$. However, this does not exclude that $p_{q}(w)$ is negative for some $w$.
IV Main result
Our main result can be formulated as the following theorem.
Theorem 1.
For the charging process of Sec. II, for $t_{1}\in(0,\tau)$ and $t_{2}=\tau$, if for $q=1/2$ the quasiprobability distribution of the work $w$ done in the time interval $t_{1},t_{2}]$ also takes negative values, i.e., $p_{\frac{1}{2}}(w)<0$ for some $w$, then there is quantum advantage in the charging process, i.e., $\tau\to 0$ as $N\to\infty$.
To show this result, we proceed in two steps. We show that if $p_{\frac{1}{2}}(w)<0$ for some $w$, then there is some skewness in the thermodynamic limit ($N\to\infty)$ and higher cumulants cannot be neglected for small deviations. Then, we show that in this case the sufficient condition of Sec. II.1 holds so that there is charging quantum advantage.
In order to derive the result, we consider the characteristic function $\chi_{q}(u)=\langle e^{iwu}\rangle$, such that the quasiprobability distribution is achieved by performing the Fourier transform $p_{q}(w)=\frac{1}{2\pi}\int e^{-iwu}\chi_{q}(u)du$. We get $\chi_{q}(u)=\frac{1}{2}(X_{q}(u)+X_{1-q}(u))$, where
$X_{q}(u)=\langle e^{-iu(1-q)H_{1}}e^{iuH_{0}}e^{-iuqH_{1}}\rangle_{\tau}\,.$ (10)
Furthermore, the cumulant generating function is defined as $G_{q}(u)=\ln\chi_{q}(u)$ and the n-th cumulant is $\kappa_{n}=(-i)^{n}G_{q}^{(n)}(0)$. For instance, for $q=1/2$ we get $\chi_{1/2}(u)=X_{1/2}(u)$. In particular, since $X_{1/2}(-u)=X^{*}_{1/2}(u)$, the Fourier transform of $X_{1/2}(u)$ takes real values, and thus $p_{1/2}(w)$ is a real valued distribution. In contrast, for $q=0$, we get $\chi_{0}(u)=X_{0}(u)=X_{1}(u)=e^{iuE^{0}_{max}}\langle e^{-iuH_{1}}\rangle_{\tau}$. Then, the cumulants $\kappa_{n}$ of the work for $q=0$ that are higher than the first cumulant $\kappa_{1}$ are equal to the cumulants $\kappa^{\prime}_{n}$ of the Hamiltonian $H_{1}$ coming from the generating function $G^{\prime}(u)=\ln\langle e^{-iuH_{1}}\rangle_{\tau}$, which are $\kappa^{\prime}_{n}=(-i)^{n}G^{\prime(n)}(0)$. In particular, the first cumulants are $\kappa^{\prime}_{1}=-\langle H_{1}\rangle_{\tau}$, $\kappa^{\prime}_{2}=\langle H_{1}^{2}\rangle_{\tau}-\langle H_{1}\rangle_{\tau}^{2}$ and $\kappa^{\prime}_{3}=-\langle H_{1}^{3}\rangle_{\tau}+3\langle H_{1}^{2}\rangle_{\tau}\langle H_{1}\rangle_{\tau}-2\langle H_{1}\rangle_{\tau}^{3}$. Since the first two moments do not depend on $q$, we get that the first two cumulants $\kappa_{n}$ are the same for any value of $q$. Thus, from the expression derived for $\chi_{0}(u)$, we get $\kappa_{1}=E^{0}_{max}-\langle H_{1}\rangle_{\tau}$ and $\kappa_{2}=\kappa_{2}^{\prime}$. In particular, the first cumulant $\kappa_{1}$ is the average work done in the interval $t_{1},t_{2}]$. In contrast, the third cumulant reads
$\kappa_{3}=\kappa^{\prime}_{3}-6q(1-q)\langle H_{1}\tilde{H}_{0}H_{1}\rangle_{\tau}\,$, (11)
where we defined the operation of inversion of the spectre of $H_{0}$ as $\tilde{H}_{0}=E^{0}_{max}-H_{0}$. If the eigenvalues of $\tilde{H}_{0}$ are equal to the eigenvalues of $H_{0}$, this operation can be realized through some unitary operator $U_{I}$ that squares to the identity, i.e., $\tilde{H}_{0}=U_{I}H_{0}U_{I}$ with $U_{I}^{2}=I$. In general, we define $g_{q}(u)$ such that
$X_{q}(u)=e^{iuE^{0}_{max}}e^{Ng_{q}(u)}\,.$ (12)
The variance of the work distribution depends on $g_{q}^{\prime\prime}(0)$, and it is nonzero if and only if $g^{\prime\prime}_{q}(0)\neq 0$. We note that $g^{\prime}_{q}(0)$ is imaginary, $g^{\prime\prime}_{q}(0)$ can be complex and $g^{\prime\prime}_{1/2}(0)$ is real. If $g_{q}(u)$ is intensive (i.e., $g_{q}(u)\sim{\it O}(1)$ as $N\to\infty$) and $g^{\prime\prime}_{q}(0)\neq 0$, then we get
$X_{q}\left(\frac{u}{N^{\frac{1}{2}}}\right)\sim e^{iuN^{-\frac{1}{2}}E^{0}_{max}+N^{\frac{1}{2}}g_{q}^{\prime}(0)u+\frac{g_{q}^{\prime\prime}(0)}{2}u^{2}+{\it O}(N^{-\frac{1}{2}})}\,$, (13)
from which we note that the skewness is ${\it O}(N^{-\frac{1}{2}})$ and tends to zero as $N\to\infty$. From Eq. (13) we get
$p_{q}\left(\frac{w}{N^{\frac{1}{2}}}\right)\sim\frac{1}{\sqrt{2\pi}}\text{Re}\left(\frac{\exp\left(\frac{(w-E^{0}_{max}+iNg_{q}^{\prime}(0))^{2}}{2Ng_{q}^{\prime\prime}(0)}\right)}{\sqrt{-g_{q}^{\prime\prime}(0)}}\right)\,$, (14)
which for $q=1/2$ reduces to the Gaussian
$p_{1/2}\left(\frac{w}{N^{\frac{1}{2}}}\right)\sim\frac{1}{\sqrt{2\pi}}\frac{\exp\left(\frac{(w-E^{0}_{max}+iNg_{1/2}^{\prime}(0))^{2}}{2Ng_{1/2}^{\prime\prime}(0)}\right)}{\sqrt{-g_{1/2}^{\prime\prime}(0)}}\,.$ (15)
Thus, $p_{q}(w)$ is positive for $q=1/2$ and also takes negative values for $q\neq 1/2$ if $g_{q}^{\prime\prime}(0)$ is not real. Therefore, for $g_{q}^{\prime\prime}(0)\neq 0$, we deduce that for $q\neq 1/2$ if $p_{q}(w)\geq 0$ (for small deviations) as $N\to\infty$, then $g_{q}(u)$ is unbounded or $g_{q}^{\prime\prime}(0)$ is real. This result is in agreement with the expression obtained for $X_{0}(u)$, which gives $p_{0}(w)\geq 0$ and $g_{0}^{\prime\prime}(0)$ real.
Furthermore, for $q=1/2$ if $p_{1/2}(w)<0$ for some $w$ as $N\to\infty$, then $g_{1/2}(u)$ is unbounded. This means that for some $n$ the derivative $g^{(n)}_{1/2}(0)$ is unbounded as $N\to\infty$. The cumulant generating function is $G_{1/2}(u)=iuE^{0}_{max}+Ng_{1/2}(u)$ and $\kappa_{n}=(-i)^{n}Ng^{(n)}_{1/2}(0)$ for $n\geq 2$. We focus on systems with charging Hamiltonian $H_{1}$ such that if $|g_{1/2}(u)|\to\infty$ as $N\to\infty$, then $|g^{\prime\prime\prime}_{1/2}(0)|\to\infty$. Thus, we deduce that if $p_{1/2}(w)<0$ for some $w$ then $\kappa_{3}/N$ is unbounded as $N\to\infty$. Of course the converse statement does not hold.
For the charging process under consideration, $\kappa_{1}\sim N$, and $\kappa_{2}=\kappa_{2}^{\prime}$. We focus on systems with charging Hamiltonian $H_{1}$ such that $\kappa^{\prime}_{3}/N$ is bounded as $N\to\infty$. For instance, if a central limit theorem holds for $H_{1}$, $\kappa_{2}^{\prime}\sim N$, then $\kappa_{n}/N$ is bounded as $N\to\infty$ for $n=1,2$. For $n\geq 3$, for $q=0$ we get $\kappa_{n}/N\sim{\it O}(1)$, however for $q\neq 0,1$ we can get $\kappa_{n}/N$ unbounded as $N\to\infty$. Let us assume that this happens for $n=3$, this can be possible due to the presence of the term $\langle H_{1}\tilde{H}_{0}H_{1}\rangle_{\tau}$ in $\kappa_{3}$. Then, for $p_{1/2}(w)$, we can get some skewness $\gamma=\kappa_{3}/\kappa_{2}^{3/2}$, e.g., if $\kappa_{3}\sim N\kappa_{2}$ we get $\gamma\sim N^{1/2}$. In the end, we note that the term $\langle H_{1}\tilde{H}_{0}H_{1}\rangle_{\tau}$ in $\kappa_{3}$, appearing for $q\neq 0,1$, scales as $\langle H_{1}H_{0}H_{1}\rangle_{0}$. In particular, we get the equality $\langle H_{1}\tilde{H}_{0}H_{1}\rangle_{\tau}=\langle H_{1}H_{0}H_{1}\rangle_{0}$ if the inversion symmetry $U_{I}H_{1}U_{I}=H_{1}$ is satisfied. Then, if $p_{1/2}(w)<0$ for some $w$, $\kappa_{3}/N$ and thus $\langle H_{1}H_{0}H_{1}\rangle_{0}/N$ are unbounded as $N\to\infty$. Then, from Sec. II.1, $\sum_{i}w_{i2}/N$ is unbounded and $\tau\to 0$, thus there is quantum advantage. In particular, if $\kappa_{3}\sim\sum_{i}w_{i2}\sim N^{2z+1}$ with an exponent $z>0$, we deduce that $\tau\sim N^{-z}$. In the end, we note that the converse of theorem 1 does not hold, since in principle we can achieve $\kappa_{3}/N$ unbounded for a nonnegative probability distribution $p_{1/2}(w)$, thus negativity gives just a sufficient condition to achieve $\tau\to 0$.
V Example
We focus on a local Hilbert space (of a single cell) having dimension $2$. Thus, we consider $h_{i}=\epsilon_{0}(\sigma^{y}_{i}+1)$, where $\sigma^{x}_{i}$, $\sigma^{y}_{i}$ and $\sigma^{z}_{i}$ are the local Pauli matrices of the i-th cell. We consider the charging Hamiltonian
$H_{1}=i\sum_{i<j}x_{ij}J_{ij}\gamma_{i}\gamma_{j}\,$, (16)
where $i,j=1,\ldots,2N$, $\{\gamma_{i},\gamma_{j}\}=2\delta_{i,j}$ and $J_{ij}\sim{\it O}(1)$ random with zero average. We consider $x_{ij}=1$ with probability $p$ and $x_{ij}=0$ with probability 1p1-p, defining the connectivity of the fermions. The average number of 2-point connections $C_{2}$ can be calculated as the average number of nonzero couplings so that $C_{2}=\sum_{i<j}\overline{x_{ij}}\sim pN^{2}$, where the bar denotes the average over the disorder. Here, for our purposes we consider the power law $C_{2}\sim N^{\alpha}$, with $0\leq\alpha\leq 2$. Let us focus on $\alpha\in[1,2]$. In this case there are ${\it O}(N)$ different fermions appearing in $H_{1}$. Then, roughly speaking, the nonzero single particle energies $\epsilon_{k}$ are such that $1\lesssim\epsilon_{k}\lesssim N$. We recall (see, e.g., Ref. Kitaev01 ) that the $N$ single particle energies $\epsilon_{k}\geq 0$ can be obtained from the eigenvalues $\pm i\epsilon_{k}$ of the $2N\times 2N$ skew-symmetric matrix $A$ with elements $A_{ij}=2x_{ij}J_{ij}$. Furthermore, we have $\sum_{k}\epsilon_{k}^{2}\sim\sum_{i<j}x_{ij}J_{ij}^{2}\sim N^{\alpha}$ since $J_{ij}\sim{\it O}(1)$. Thus, we aim to look for a disorder $J_{ij}\sim{\it O}(1)$ such that $\Delta E_{1}=\sum_{k}\epsilon_{k}\sim N$. This is possible if we have only $N^{c}$ energies $\epsilon_{k}\neq 0$ such that $\epsilon_{k}\sim N^{d}$. In this case we get $\sum_{k}\epsilon_{k}^{2}\sim N^{c+2d}$ and $\sum_{k}\epsilon_{k}\sim N^{c+d}$, from which we deduce the solution $c=2-\alpha$ and $d=\alpha-1$. Let us focus on $\alpha\in[0,1]$. In this case only ${\it O}(N^{\alpha})$ fermions appear in the Hamiltonian $H_{1}$, and thus $\Delta E_{1}\sim N^{\alpha}$. Then, for $\alpha\in[0,1]$ in order to get a gap $\Delta E_{1}\sim N$, we consider
$H_{1}=iN^{1-\alpha}\sum_{i<j}x_{ij}J_{ij}\gamma_{i}\gamma_{j}\,$, (17)
which has ${\it O}(N^{\alpha})$ energies $\epsilon_{k}\neq 0$ which scale as $\epsilon_{k}\sim N^{1-\alpha}$. In particular, the Hamiltonian in Eq. (16) can be achieved from the Hamiltonian in Eq. (17) by performing a linear transformation, i.e., $\gamma_{i}\mapsto\sum_{j}W_{ij}\gamma_{j}$, where $W_{ij}$ is an orthogonal matrix. In this terms, the Hamiltonian in Eq. (17) with $\alpha=\alpha^{\prime}$ is the diagonal form of the Hamiltonian in Eq. (16) with $\alpha=2-\alpha^{\prime}$. Thus, the disorder $J_{ij}\sim{\it O}(1)$ for $\alpha=2$ can be realized as $J_{kl}=N\sum_{i,j}W_{ki}x_{ij}J^{\prime}_{ij}W_{jl}\sim{\it O}(1)$ (which is correlated) where $J_{ij}=J^{\prime}_{ij}\sim{\it O}(1)$ is the random coupling with zero average of the Hamiltonian in Eq. (17) with $\alpha=0$. Thus, we write
$H_{1}=iM\sum_{i<j}x_{ij}J_{ij}\gamma_{i}\gamma_{j}\,$, (18)
where $M=1$ for $\alpha\in[1,2]$ and $M=N^{1-\alpha}$ for $\alpha\in[0,1]$. In detail, the operators $\gamma_{i}$ are related to the Pauli matrices $\sigma^{x}_{i}$, $\sigma^{y}_{i}$ and $\sigma^{z}_{i}$ through the Jordan-Wigner transformation
$\gamma_{2l}$ $=$ $\left(\prod_{k=1}^{l-1}\sigma^{z}_{k}\right)\sigma^{x}_{l}\,$, (19)
$\gamma_{2l-1}$ $=$ $\left(\prod_{k=1}^{l-1}\sigma^{z}_{k}\right)\sigma^{y}_{l}\,.$ (20)
Then, we deduce the inversion operator $U_{I}=\otimes_{i}\sigma^{z}_{i}$, which is such that $U_{I}H_{0}U_{I}=\sum_{i}\tilde{h}_{i}$, with $\tilde{h}_{i}=\epsilon_{0}(1-\sigma^{y}_{i})$. Furthermore, we get $U_{I}H_{1}U_{I}=H_{1}$ since $U_{I}$ is the parity operator. We assume that $|\psi(\tau)\rangle\approx|E^{0}_{max}\rangle$, so that the main result applies at least approximately. Since $\langle H_{1}h_{i}H_{1}\rangle_{0}\sim\Delta H_{1}^{2}$, where $\Delta H_{1}^{2}=\kappa^{\prime}_{2}$, for a number $N^{\prime}\sim N$ of indices $i$, we get a quantum advantage with $\tau\sim N^{-z}$ and exponent $z=1-\alpha/2$ for $\alpha\in[0,1]$ and $z=\alpha/2$ for $\alpha\in[1,2]$ (see Ref. Francica25 ). Thus, for $q=1/2$, we get $\kappa_{1}\sim N$, $\kappa_{2}\sim N^{2z}$ and $\kappa_{3}\sim N^{2z+1}$. To get some insights concerning the negativity of $p_{1/2}(w)$, from the series $G_{q}(u)=\sum_{n}i^{n}\kappa_{n}u^{n}/n!$ by performing a truncation we get the result $G_{1/2}(u)\approx i\kappa_{1}u-\kappa_{2}u^{2}/2-i\kappa_{3}u^{3}/6$. If we neglect $\kappa_{2}$, by performing a Fourier transform for $\kappa_{3}<0$ we get
$p_{1/2}(w)\sim Ai\left(\sqrt[3]{\frac{2}{|\kappa_{3}|}}(\kappa_{1}-w)\right)$ (21)
and the Airy function shows oscillations and takes negative values for some $w>\kappa_{1}$. However, the truncation leads to a result that can be far from the exact one. For this model we can try to calculate $X_{q}(u)$ as follows. We note that
$H_{1}^{2}=M^{2}\sum_{i<j}x_{ij}J_{ij}^{2}I+\cdots$ (22)
so that we define $J_{1}=\sum_{i<j}x_{ij}J_{ij}^{2}\sim N^{\alpha}$, where we are omitting some irrelevant terms. We get
$H_{1}^{2l}=M^{2l}(J_{1})^{l}I+\cdots$ (23)
and, by noting that $H_{1}^{2l-1}=H_{1}^{2l-2}H_{1}$,
$H_{1}^{2l-1}=M^{2l-2}(J_{1})^{l-1}H_{1}+\cdots$ (24)
for $l=1,2,\ldots$. From an expansion in series of the exponentials with $H_{1}$ we get
$X_{q}(u)$ $=$ $C_{q}(u)C_{1-q}(u)e^{iuE^{0}_{max}}-S_{q}(u)S_{1-q}(u)$ (25)
$\times\langle H_{1}e^{iuH_{0}}H_{1}\rangle_{\tau}+\cdots$
where we noted that $|\psi(\tau)\rangle\approx|E^{0}_{max}\rangle$ and $\langle\gamma_{i}\gamma_{j}\rangle_{\tau}=0$ for $i\neq j$. The functions $C_{q}(u)$ and $S_{q}(u)$ read
$C_{q}(u)=\cos(uqM\sqrt{J_{1}})\,,\quad S_{q}(u)=\frac{\sin(uqM\sqrt{J_{1}})}{M\sqrt{J_{1}}}\,.$ (26)
Thus, by noting that
$\langle H_{1}e^{iuH_{0}}H_{1}\rangle_{\tau}\approx e^{iuE^{0}_{max}}\langle H_{1}e^{-iuH_{0}}H_{1}\rangle_{0}$ (27)
due to the symmetry $U_{I}$, it is easy to see that
$X_{q}(u)$ $\approx$ $C_{1}(u)e^{iuE^{0}_{max}}+S_{q}(u)S_{1-q}(u)e^{iuE^{0}_{max}}$ (28)
$\times\left(M^{2}J_{1}-\langle H_{1}e^{-iuH_{0}}H_{1}\rangle_{0}\right)+\cdots\,.$
For $q=0$, we get $X_{0}(u)=C_{1}(u)e^{iuE^{0}_{max}}+\cdots$, so that the Fourier transform is positive, i.e., $p_{0}(w)\geq 0$. Thus, $X_{q}(-u)\approx X_{q}^{*}(u)$, then its Fourier transform is real, and $X_{q}(u)\approx X_{1-q}(u)$, then $\chi_{q}(u)\approx X_{q}(u)$. From Eq. (28) we get the first cumulants $\kappa_{1}\approx E^{0}_{max}$, $\kappa_{2}\approx M^{2}J_{1}$ and $\kappa_{3}\approx-6q(1-q)\langle H_{1}H_{0}H_{1}\rangle_{0}$, while $\kappa^{\prime}_{1}=0$, $\kappa^{\prime}_{2}\approx\kappa_{2}$ and $\kappa^{\prime}_{3}\approx 0$. Thus, $\kappa_{3}\sim N^{2z+1}$ for $q\neq 0,1$ and the distribution of work in this case can be negative. Clearly, the negativity for $q\neq 0,1$ comes from the second term in Eq. (28), which is zero for $u=0$. In general, the Fourier transform of a function $S(u)$, such that $S(0)=0$, is $s(w)=\frac{1}{2\pi}\int e^{-iwu}S(u)du$, so that $S(0)=\int s(w)dw=0$ implies that $s(w)$ takes both positive and negative values or $s(w)=0$ for all $w$. The exact quasiprobability distribution can be numerically calculated for not too large $N$, e.g., for $\alpha=1$ we see how $p_{0}(w)$ tends to be a Gaussian and $p_{1/2}(w)$ also takes negative values for some $w<\kappa_{1}$ (see Fig. 1).
Figure 1: The histograms of quasiprobability distributions $p_{0}(w)$ (top panel) and $p_{1/2}(w)$ (bottom panel) for a single realization of the disorder. We assume $|\psi(\tau)\rangle\approx|E^{0}_{max}\rangle$ and put $C_{2}=N^{\alpha}$ with $\alpha=1$, $N=8$ and $\epsilon_{0}=1$.
VI Conclusions
We showed how the quantum advantage is connected to the quasiprobability of the work done in a subinterval of the charging process. In particular, from the point of view of the work statistics, the work always admits a representation that is a (non-contextual) probability distribution. We hope that this can help us to better understand the origin of the quantum advantage. Thus, if we turn on a charging interaction for a certain duration time with the aim of charging the battery, depending on the duration time the battery can be charged or not. If there is quantum advantage, the expected time that is needed to charge the battery tends to zero as the number of cells increases. Beyond theoretical purposes, this suggests that our result can find some applications in the realization of quantum batteries. For example, if the quantum advantage is not achieved correctly (something went wrong with the charging process), the battery will not be charged in the expected duration time. In order to check if the battery will be charged in that duration time, we can look at the energy of the battery. Of course, the battery will be charged if the quantum advantage is achieved. Then, if we look on the negativity of a quasiprobability distribution, the battery will surely be charged in the expected time if the quasiprobability also takes negative values (because in this case quantum advantage will be achieved). Thus, we hope that our results can be useful for the realization of charging setups, since the quasiprobability distribution, and thus its negativity, can be measured by coupling a detector to the charging setup, without destroying the state of the battery (see, e.g., Ref. Francica22a ). Basically, to answer the question “is the battery charged in the (quantum advantage) expected duration time?” all the necessary information is yes or no, and can be obtained by answering the question “there is negativity?”.